OpenStudy (anonymous):

There are 3 taps for three pipes x, y and z. x and y are inlet pipes, which can fill the tank in 20 hours and 25 hours respectively, whereas z is an outlet tap. When the tank is empty, a man want to turn on x and y, but by mistake he turned on x and z. It takes 10 hours more to fill the tank than expected. In how much time can pipe z empty the tank alone

OpenStudy (australopithecus):

This looks like a simple system of equations problem. Write out formulas and do system of equations to figure out z

OpenStudy (anonymous):

1/20+1/25=1/z

OpenStudy (anonymous):

1/20+9/100=10

OpenStudy (aum):

If x and y are turned on, how long will it take to fill the tank? Use 1/x + 1/y = 1/t . t = xy / (x+y) If x and z are turned on, how long will it take to fill the tank? Use 1/x - 1/z = 1/t1 t1 = xz / (z-x) "It takes 10 hours more to fill the tank than expected." t1 - t = 10 xz / (z-x) - xy / (x+y) = 10 Put x = 20, y = 25 and solve for z.

OpenStudy (anonymous):

z=380